# Mosaic-IT: Free Compositional Data Augmentation Improves Instruction Tuning

### v2

@@ Abstract
Finetuning large language models with a variety of instruction-response pairs has enhanced their capability to understand and follow instructions. Current instruction tuning primarily relies on teacher models or human intervention to generate and refine the instructions and responses for training, which are costly, non-sustainable, and may lack diversity. In this paper, we introduce Mosaic Instruction Tuning (Mosaic-IT), a human/model-free compositional data augmentation method that can efficiently create rich and diverse augmentations from existing instruction tuning data to enhance the LLMs. Mosaic-IT randomly concatenates multiple instruction data into one and trains the model to produce the corresponding responses with predefined higher-level meta-instructions to strengthen its multi-step instruction-following and format-following skills. Our extensive evaluations demonstrate a superior performance and training efficiency of Mosaic-IT, which achieves consistent performance improvements over various benchmarks and a $80\%$ reduction in training costs compared with original instruction tuning. Our codes and data are available at https://github.com/tianyi-lab/Mosaic-IT.
@@ 1 Introduction
The emergence of Large Language Models (LLMs) Brown et al. (2020); Scao et al. (2022); OpenAI (2023); Touvron et al. (2023a; b); Jiang et al. (2023) along with their remarkable performance in down-stream tasks Zhao et al. (2023); Xu et al. (2024a), has revolutionized the domains of Artificial Intelligence and Natural Language Processing. A key component of the recipe to unlock the exceptional ability of LLMs in understanding and following instructions is the technique of Instruction Tuning (IT) Mishra et al. (2021); Wei et al. (2022); Chung et al. (2022); Wang et al. (2023c); Zhang et al. (2023); Xu et al. (2024a), which involves the fine-tuning of LLMs on datasets comprising corresponding instruction-response pairs.
To ensure the quality of instruction tuning data, earlier efforts Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a); Jiang et al. (2023) carefully curate extensive, diverse, and high-quality datasets manually. Although these datasets encompass a wide range of instructions to improve instruction tuning, they require the responses to be meticulously curated by human experts Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022). Alternatively, some approaches Wang et al. (2023b); Taori et al. (2023); Xu et al. (2023); Li et al. (2023a) leverage more capable teacher LLMs to reduce the labor-intensive process of data generation. For example, the Alpaca Taori et al. (2023) utilizes self-instruct Wang et al. (2023b) to automatically generate diverse instruction tuning datasets, and the WizardLM Xu et al. (2023) proposes to complicate the existing instruction data by an evolution algorithm. Building on this trend and the widely acknowledged notion that more complicated instructions are more beneficial for LLMs’ instruction-following ability Xu et al. (2023); Zhao et al. (2024), numerous strategies Zhao et al. (2024); Wu et al. (2024); Ding et al. (2023); Li et al. (2023a); Liu et al. (2023a); Li et al. (2024b; a); Guo et al. (2024); Xu et al. (2024a) have been proposed to further diversify and complexify the instruction-response pairs, utilizing teacher models like ChatGPT-3.5 and GPT-4 OpenAI (2023).
Despite the enhanced performance in instruction-following ability offered by these existing methods, they face Two major issues: (1) They heavily rely on teacher models or human annotators to rewrite instruction-response pairs, which highlights the resource-intensive nature and their constraints on scalability; (2) They only increase the complexity within the scope of a single instruction, which limits the potential improvement in LLMs’ instruction-following capabilities. Motivated by the Dense and Aligned Captions Doveh et al. (2023) proposed for vision language (VL) models and the mosaic data augmentation proposed in Yolov4 Bochkovskiy et al. (2020), we hypothesize that denser instructions benefit the LLM alignment, i.e. the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. A similar concept during the inference phase is proposed by batch prompting Cheng et al. (2023); Lin et al. (2024), where multiple samples are grouped in one batch allowing LLMs to generate multiple responses at one inference, while its performances are sub-optimal. Moreover, our preliminary experiments on GPT-3.5-turbo and GPT-4-turbo show that even for these strong proprietary LLMs, their performances degrade dramatically if required to follow several instructions at one time, the experimental results are presented in the Section 5 Further Discussion. Thus, these performance degradation phenomenons indicate the complexity of this setting and the necessity of further training for this higher-level capability.
@@
Figure 1: The illustration of our Mosaic-IT with different strategies. Given the original dataset, our method randomly samples and concatenates them together into more complex samples, simulating the multi-instruction-following scenarios at no cost.
@@
Though effective, the Primary Mosaic strategy constrains LLMs in responding to the instructions in the original order and format, potentially limiting its further potential. Thus we further introduce three Advanced Mosaic Strategies aimed at enhancing the diversity and complexity of the mosaicked instruction-response pairs: Format, Permute, and Maskout, in which an additional meta-instruction is provided as a higher-level guideline for LLMs to follow the given instructions. Illustrative examples are presented in Figure 2. Specifically, in the Format strategy, some arbitrary parsing formats will be defined in the meta-instruction thus forcing LLMs to follow these formats, which notably enhances the LLMs’ capacity to follow formats. In the Permutation strategy, an arbitrary permuted order is defined thus forcing LLMs to respond in a desired order. In the Maskout strategy, some arbitrary instructions are sampled which meta-instruction forces LLMs to ignore. Moreover, the use of these Advanced strategies not only boosts the performance in several evaluation metrics but also keeps our method free of additional costs.
@@
  • We propose a novel human/model-free data augmentation method, Mosaic-IT, which extends existing instruction tuning from handling one single instruction at a time to following multiple instructions in diverse forms. This approach significantly enhances the potential utilization of existing high-quality datasets.
@@
Figure 2: Illustrative examples of Mosaic-IT. Given $3$ simple data points, our method can concatenate them into overall data samples with diverse forms. Texts in red represent the meta-instructions that define the formats or orders for LLMs to respond. Texts in yellow are major response differences of each strategy. The Primary Strategy only concatenates data together. The Format Strategy requires LLMs to respond in predefined formats. The Permute Strategy requires LLMs to respond in specific orders and the Maskout Strategy requires LLMs to ignore some of the instructions.
@@ 2.1 Preliminaries
For simplicity, we define $x=\texttt{map}(\textit{Instruction},\textit{Input})$ as the unified instruction, and $y$ as the corresponding response. Therefore, $D$ can be represented as ${(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{n},y_{n})}$, denoting a set of $n$ instruction-response pairs. Let $p_{\theta}(\cdot)$ denote the LLMs to be trained, with parameters $\theta$. In the instruction tuning setting, $p_{\theta}$ is typically fine-tuned by maximizing the following objective on each data $(x_{i},y_{i})$, $y_{i,j}$ represents the $j_{th}$ token of response $y_{i}$, $y_{i,<j}$ represents the tokens prior to $y_{i,j}$, and $l_{i}$ represents the token length of $y_{i,j}$:
@@ 2.2 Mosaic-IT
Motivated by the success of the existing data-centric instruction tuning methods, a line of approaches is proposed to further enhance the instruction-response pairs utilizing extra teacher LLMs Xu et al. (2024a). Though effective, all existing methods for instruction tuning restrict training samples to just one instruction, which severely limits the potential of the existing high-quality data and the instruction-following ability of the models to be trained. Motivated by the Dense and Aligned Captions Doveh et al. (2023) for VL, we hypothesize that denser instructions benefit the LLM alignment, thus the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. Thus, we propose Mosaic Instruction Tuning (Mosaic-IT) as shown in Figure 1.
@@ 2.2.1 Primary Mosaic strategy
Exploring the concept of concatenating random instruction-response pairs into a unified instruction-response pair for training remains largely unexplored. The primary challenge lies in crafting a coherent overall instruction and obtaining its corresponding response. Most existing methods utilize a strong teacher model to rewrite and polish the instructions with prompting techniques and generate corresponding responses, introducing more cost by actually re-generating new data samples. To harness the full potential of existing data rather than directly discarding them, we introduce a simple compositional approach as shown in Figure 2, in which instructions are randomly concatenated with serial digits to form an overall instruction. The concatenated overall instruction is denoted as $[x_{1},...,x_{k}]$, with the corresponding overall response concatenated as $[y_{1},...,y_{k}]$. Here, $k$ denotes the number of original data samples integrated into each overall sample.
@@ 2.2.2 Advanced Mosaic Strategies
Though effective, this simple primary mosaic strategy constrains LLMs in responding to the instructions with the original order and format, potentially limiting its generalization and practical usage. In our method, the instructions and corresponding responses from the original dataset can be viewed as atomic components and our method randomly combines these elements together to form new instructions and responses. This nature allows us to further complicate this process with fancier strategies thus forcing LLMs to follow more complicated overall instructions. Hence, we propose three Advanced Mosaic Strategies to complicate and diversify the mosaicked samples as shown in Figure 2, including Format, Permute, and Maskout, with meta-instructions guiding them.
Format In the Format strategy, some arbitrary formats are defined in the meta-instruction to force LLMs to follow these formats in the response. The formats mainly contain two categories: 1) Serial Digit Format and (2) Response Parsing Format. The serial digits establish the initial instruction order that guides LLMs to follow sequentially. We manually define $10$ types of serial digit format, which will be randomly sampled during each mosaic process. For response parsing, we simulate the scenario where the users try to extract specific information from the responses. We define $27$ types of parsing brackets and $17$ types of parsing text pairs, which will be randomly sampled and assembled during each mosaic process. Examples can be found in Appendix D, which can be easily extended for customized training settings. We denote responses with specific formats as $y^{\prime}_{i}=wrap(y_{i},s_{format})$, and $l$ as the token length of the overall response. An additional meta-instruction $s_{format}$ specifying the required format will be included in the overall instruction. Thus, the objective function for each mosaic data point:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{k}]_{j}|[x_{1},...,x_{k},s_{format}],[y^{\prime}_{1},...,y^{\prime}_{k}]_{<j}\right)$ (3)
@@
In the Permute strategy, an arbitrary permuted order is defined in the meta-instructions, forcing LLMs to follow. Moreover, several high-level rules are defined to ensure the complexity and diversity of meta-instructions, e.g., forcing LLMs to respond to each instruction in the randomly generated permutation list, forcing LLMs to respond in the alphabetical order of each instruction, forcing LLMs to respond according to the length of instructions, etc. The detailed rule types and descriptions are depicted in Appendix D. These various meta-instructions not only provide higher-level guidelines for LLMs to follow multiple instructions but also inherently enhance the instruction perception ability of LLMs. In our settings, LLMs are required to generate responses selectively conditioned on some critical parts of the overall instruction, forcing them to first understand the formats and other requirements, indicating a more comprehensive understanding of the context given. The meta-instruction is denoted as $s_{permute}$ and is included in the overall instruction. The permuted response list is denoted as $[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]=Permute([y^{\prime}_{1},...,y^{\prime}_{k}],s_{permute})$. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{j}|[x_{1},...,x_{k},s_{format},s_{permute},[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{<j}]\right),$ (4)
In the Maskout strategy, some arbitrary instructions are selected in the meta-instructions forcing LLMs to ignore them. Several high-level rules are also defined similarly to the permute strategy, including forcing LLMs to ignore the instructions with given random digits, forcing LLMs to ignore the longest one/several instructions, forcing LLMs to ignore odd-numbered instructions, etc. The details are provided in Appendix D. Similarly, the meta-instruction is denoted as $s_{maskout}$ and the response list is denoted as $[y^{\prime}_{1},...,y^{\prime}_{m}]=Maskout([y^{\prime}_{1},...,y^{\prime}_{k}],s_{maskout})$, where $m$ is the count of responses after masking out. Thus the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\left([y^{\prime}_{1},...,y^{\prime}_{m}]_{j}|[x_{1},...,x_{k},s_{format},s_{maskout}],[y^{\prime}_{1},...,y^{\prime}_{m}]_{<j}\right)$ (5)
@@
How to decide the Number of Instructions $k$: Number of Instructions denotes the number of original data samples that are integrated into an overall sample. In addition to the detailed mosaic strategies being used, this count also dramatically affects the effect of Mosaic-IT. Our experiments reveal that larger and more diverse numbers of instructions will benefit LLM training. By default, we set the maximum number of instructions as $k_{max}=10$, and randomly sample an integer that is smaller or equal to $k_{max}$ under a uniform distribution. If the number causes the data sample to be longer than the max length, it will be automatically reduced to the max number which remains the sample length within the limits.
3 Experimental Setup
3.1 Implementation Details
The experiments are conducted on: Llama2-7B, Llama2-13B Touvron et al. (2023b), and Mistral-7B Jiang et al. (2023), Llama-3-8B Dubey et al. (2024), Phi-3 Abdin et al. (2024), and Gemma2-2B Team et al. (2024). The training datasets include Alpaca Taori et al. (2023), Alpaca-GPT4 Peng et al. (2023), WizardLM Xu et al. (2023), Vicuna 1M Zheng et al. (2024a), and Magpie Xu et al. (2024b) datasets. Due to the extremely large size of Vicuna 1M and Magpie, $300$k instances are randomly sampled for our experiments. The detailed description of datasets and the training configurations are introduced in Appendix B.
3.2 Evaluation Metrics
We utilize five automatic evaluation metrics, including (i) LLM-based Pair-wise Comparison, (ii) Open LLM leaderboard, (iii) MT-Bench, (iv) Alpaca Eval, and (v) IF Eval, and (vi) Human evaluation to verify the effectiveness of our method. They are widely accepted evaluation metrics for measuring LLMs’ instruction-following capabilities. The introductions of the five automatic evaluation metrics are provided in Appendix B.

### v3

# Mosaic-IT: Cost-Free Compositional Data Synthesis for Instruction Tuning

@@ Abstract
Finetuning large language models with a variety of instruction-response pairs has enhanced their capability to understand and follow instructions. Current instruction tuning primarily relies on teacher models or human intervention to generate and refine the instructions and responses for training, which are costly, non-sustainable, and may lack diversity. In this paper, we introduce Mosaic Instruction Tuning (Mosaic-IT), a human/model-free compositional data synthesis method that can efficiently create rich and diverse augmentations from existing instruction tuning data to enhance the LLMs. Mosaic-IT randomly concatenates multiple instruction data into one and trains the model to produce the corresponding responses with predefined higher-level meta-instructions to strengthen its multi-step instruction-following and format-following skills. Our extensive evaluations demonstrate a superior performance and training efficiency of Mosaic-IT, which achieves consistent performance improvements over various benchmarks and an $80\%$ reduction in training costs compared with original instruction tuning. Our codes and data are available at https://github.com/tianyi-lab/Mosaic-IT.
Mosaic-IT: Cost-Free Compositional Data Synthesis for Instruction Tuning
Ming Li1, Pei Chen2, Chenguang Wang3, Hongyu Zhao1, Yijun Liang1 Yupeng Hou2, Fuxiao Liu1, Tianyi Zhou1 1University of Maryland     2Texas A&M University     3Stony Brook University minglii@umd.edu    tianyi.david.zhou@gmail.com
@@ 1 Introduction
The emergence of Large Language Models (LLMs) Scao et al. (2022); OpenAI (2023); Touvron et al. (2023a) along with their remarkable performance in downstream tasks Zhao et al. (2023); Xu et al. (2024a), has revolutionized the domains of Artificial Intelligence and Natural Language Processing. A key component of the recipe to unlock the exceptional ability of LLMs in understanding and following instructions is the technique of Instruction Tuning (IT) Mishra et al. (2021); Wei et al. (2022); Chung et al. (2022), which involves the fine-tuning of LLMs on datasets comprising corresponding instruction-response pairs.
@@
Figure 1: The brief illustration of our Mosaic-IT with different strategies. Given the original dataset, our method randomly samples and concatenates them together into more complex samples, simulating the multi-instruction-following scenarios in a cost-free manner.
To ensure the quality of instruction tuning data, earlier efforts Brown et al. (2020); OpenAI (2023); Touvron et al. (2023a); Jiang et al. (2023) carefully curate extensive, diverse, and high-quality datasets manually. Although these datasets encompass a wide range of instructions to improve instruction tuning, they require the responses to be meticulously curated by human experts Khashabi et al. (2020); Ye et al. (2021); Wei et al. (2022); Wang et al. (2022); Du et al. (2022). Alternatively, some approaches Wang et al. (2023b); Taori et al. (2023); Xu et al. (2023); Li et al. (2023a) leverage more capable teacher LLMs to reduce the labor-intensive process of data generation. For example, the Alpaca Taori et al. (2023) utilizes self-instruct Wang et al. (2023b) to automatically generate diverse instruction tuning datasets. Building on this trend and the widely acknowledged notion that more complicated instructions are more beneficial for LLMs’ instruction-following ability Xu et al. (2023); Zhao et al. (2024), numerous strategies Zhao et al. (2024); Wu et al. (2024); Ding et al. (2023); Li et al. (2023a); Liu et al. (2023a); Li et al. (2024c, b); Guo et al. (2024); Xu et al. (2024a) have been proposed to further diversify and complexity the instruction-response pairs, utilizing teacher models like ChatGPT-3.5 and GPT-4 OpenAI (2023).
Despite the enhanced performance in instruction-following ability offered by these existing methods, they face Two major issues: (1) They heavily rely on teacher models or human annotators to rewrite instruction-response pairs, which highlights the resource-intensive nature and their constraints on scalability; (2) They only increase the complexity within the scope of a single instruction, which limits the potential improvement in LLMs’ instruction-following capabilities. Motivated by the Dense and Aligned Captions Doveh et al. (2023) proposed for vision language (VL) models and the mosaic data augmentation proposed in Yolov4 Bochkovskiy et al. (2020), we hypothesize that denser instructions benefit the LLM alignment, i.e. the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. A similar concept during the inference phase is proposed by batch prompting Cheng et al. (2023); Lin et al. (2024), where multiple samples are grouped in one batch, allowing LLMs to generate multiple responses at one inference, while its performance is sub-optimal.
@@
Though effective, the Primary Mosaic strategy constrains LLMs in responding to the instructions in the original order and format, potentially limiting their further potential. Thus, we further introduce three Advanced Mosaic Strategies aimed at enhancing the diversity and complexity of the mosaicked instruction-response pairs: Format, Permute, and Maskout, in which an additional meta-instruction is provided as a higher-level guideline for LLMs to follow the given instructions. Illustrative examples are presented in Figure 2. Specifically, in the Format strategy, some arbitrary parsing formats will be defined in the meta-instruction, thus forcing LLMs to follow these formats, which notably enhances the LLMs’ capacity to follow formats. In the Permutation strategy, an arbitrary permuted order is defined, thus forcing LLMs to respond in a desired order. In the Maskout strategy, some arbitrary instructions are sampled, which meta-instruction forces LLMs to ignore. Moreover, the use of these Advanced strategies not only boosts the performance in several evaluation metrics but also keeps our method free of additional costs.
@@
  • We propose the first cost-free data Synthesis method, Mosaic-IT, which extends existing instruction tuning from handling one single instruction at a time to following multiple instructions in diverse forms. This approach significantly enhances the potential utilization of existing high-quality datasets.
@@
Figure 2: Illustrative examples of Mosaic-IT. Given $3$ simple data instances, our method concatenates them into data samples with diverse forms. Texts in red represent the meta-instructions that define the formats or orders for LLMs to respond. Texts in yellow are major response differences of each strategy. The Primary Strategy only concatenates data. The Format Strategy requires LLMs to respond in predefined formats. The Permute Strategy requires LLMs to respond in specific orders, and the Maskout Strategy requires LLMs to ignore some instructions.
@@ 2.2 Mosaic-IT
Motivated by the success of the existing data-centric instruction tuning methods, a line of approaches is proposed to further enhance the instruction-response pairs utilizing extra teacher LLMs Xu et al. (2024a). Though effective, all existing methods for instruction tuning restrict training samples to just one instruction, which severely limits the potential of the existing high-quality data and the instruction-following ability of the models to be trained. Motivated by the Dense and Aligned Captions Doveh et al. (2023) for VL, we hypothesize that denser instructions benefit the LLM alignment, thus the process of instruction tuning should not be constrained by one single instruction but be extended to follow several instructions at a time, which represents a higher level of instruction-following ability that is beneficial to the training process. Thus, we propose the cost-free data synthesis method, Mosaic Instruction Tuning (Mosaic-IT), as shown in Figure 1.
@@ 2.2.1 Primary Mosaic strategy
Exploring the concept of concatenating random instruction-response pairs into a unified instruction-response pair for training remains largely unexplored. The primary challenge lies in crafting a coherent overall instruction and obtaining its corresponding response. Most existing methods utilize a strong teacher model to rewrite the instructions with prompting techniques and generate corresponding responses, introducing more cost by actually re-generating new data samples. To harness the full potential of existing data rather than directly discarding it, we introduce a simple cost-free compositional approach as shown in Figure 2, in which instructions are randomly concatenated with serial digits to form an overall instruction. The concatenated overall instruction is denoted as $[x_{1},...,x_{k}]$, with the corresponding overall response concatenated as $[y_{1},...,y_{k}]$. Here, $k$ denotes the number of original data samples integrated into each overall sample.
@@ 2.2.2 Advanced Mosaic Strategies
Though effective, this primary mosaic strategy constrains LLMs in responding to the instructions with the original order and format, potentially limiting their generalization and practical usage. In our method, the instructions and corresponding responses from the original dataset can be viewed as atomic components, and our method randomly combines these elements to form new instructions and responses. This nature allows us to further complicate this process with fancier strategies, thus forcing LLMs to follow more complicated overall instructions. Hence, we propose three Advanced Mosaic Strategies to complicate and diversify the mosaicked samples as shown in Figure 2, including Format, Permute, and Maskout, with meta-instructions guiding them. These strategies are still purely rule-based Li et al. (2024a) and do not incorporate the additional human/LLM generation.
Format In the Format strategy, some arbitrary formats are defined in the meta-instruction to force LLMs to follow these formats in the response. The formats mainly contain two categories: 1) Serial Digit Format and (2) Response Parsing Format. The serial digits establish the initial instruction order that guides LLMs to follow sequentially. We manually define $10$ types of serial digit format, which will be randomly sampled during each mosaic process. For response parsing, we simulate the scenario where the users try to extract specific information from the responses. We define $27$ types of parsing brackets and $17$ types of parsing text pairs, which will be randomly sampled and assembled during each mosaic process. Examples can be found in Appendix G, which can be easily extended for customized training settings. We denote responses with specific formats as $y^{\prime}_{i}=wrap(y_{i},s_{format})$, and $l$ as the token length of the overall response. An additional meta-instruction $s_{format}$ specifying the required format will be included in the overall instruction. Thus, the objective function for each mosaic data point:
$\max_{\theta}\sum_{j=1}^{l}\log p_{\theta}\Big{(}$ $[y^{\prime}_{1},...,y^{\prime}_{k}]_{j}\Big{|}[x_{1},$ (3)
$x_{2},...,x_{k},s_{format}],[y^{\prime}_{1},...,y^{\prime}_{k}]_{<j}\Big{)}$
@@
In the Permute strategy, an arbitrary permuted order is defined in the meta-instructions, forcing LLMs to follow. Moreover, several high-level rules are defined to ensure the complexity and diversity of meta-instructions, e.g., forcing LLMs to respond to each instruction in the randomly generated permutation list, forcing LLMs to respond in the alphabetical order of each instruction, forcing LLMs to respond according to the length of instructions, etc. The detailed rule types and descriptions are depicted in Appendix G. These various meta-instructions not only provide higher-level guidelines for LLMs to follow multiple instructions but also inherently enhance the instruction perception ability of LLMs. In our settings, LLMs are required to generate responses selectively conditioned on some critical parts of the overall instruction, forcing them to first understand the formats and other requirements, indicating a more comprehensive understanding of the context given. The meta-instruction is denoted as $s_{permute}$ and is included in the overall instruction. The permuted response list is denoted as $[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]=Permute([y^{\prime}_{1},...,y^{\prime}_{k}],s_{permute})$. Thus, the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}$ $\log p_{\theta}\Big{(}[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{j}\Big{|}[x_{1},$ (4)
$x_{2},...,x_{k},s_{format},s_{permute}],[y^{\prime}_{1^{\prime}},...,y^{\prime}_{k^{\prime}}]_{<j}\Big{)}$
In the Maskout strategy, some arbitrary instructions are selected in the meta-instructions, forcing LLMs to ignore them. Several high-level rules are also defined similarly to the permute strategy, including forcing LLMs to ignore the instructions with given random digits, forcing LLMs to ignore the longest one/several instructions, forcing LLMs to ignore odd-numbered instructions, etc. The details are provided in Appendix G. Similarly, the meta-instruction is denoted as $s_{maskout}$ and the response list is denoted as $[y^{\prime}_{1},...,y^{\prime}_{m}]=Maskout([y^{\prime}_{1},...,y^{\prime}_{k}],s_{maskout})$, where $m$ is the count of responses after masking out. Thus, the objective function can be formulated as below:
$\max_{\theta}\sum_{j=1}^{l}$ $\log p_{\theta}\Big{(}[y^{\prime}_{1},...,y^{\prime}_{m}]_{j}\Big{|}[x_{1},$ (5)
$x_{2},...,x_{k},s_{format},s_{maskout}],[y^{\prime}_{1},...,y^{\prime}_{m}]_{<j}\Big{)}$
@@
How to decide the Number of Instructions $k$: Number of Instructions denotes the number of original data samples that are integrated into an overall sample. In addition to the detailed mosaic strategies being used, this count also dramatically affects the effect of Mosaic-IT. Our experiments reveal that larger and more diverse numbers of instructions will benefit LLM training. By default, we set the maximum number of instructions as $k_{max}=10$, and randomly sample an integer that is smaller than or equal to $k_{max}$ under a uniform distribution. If the number causes the data sample to be longer than the max length, it will be automatically reduced to the max number, which remains the sample length within the limits.